Smith Company manufactures fish tanks. One of the fish tanks that the company produces is a 2.6 gallon fish tank that the conpany sells foe $37.50. The fish tanks are manufactured in an outdated fish tank manufacturing facility that is labor intense, relying heavily on direct labor workers. Variable costs are high, totaling $22.50 per 2.6 gallon fish tank of whick 60% is direct labor cost. Last year the company sold 40,000 2.6 gallon fish tanks and had $480,000 in fixed expense associated with the 2.6 gallon fish tank. Compute the contribution margin ratio, break-even in number of 2.6 gallon fish tanks, and the degree of operating leverage at last year's level of sales.
Answer #1

Answer

--All working forms part of the answer

A Sale price per unit $37.50
B Variable cost per unit $22.50
C = A - B Contribution margin per unit $15.00
D = (C/A) x 100 Contribution margin ratio 40% Answer #1
E Fixed Cost $480,000
F = E/C Break even point in numbers 32000 Answer #2
G Actual unit sales 40000
H = G x C Total Contribution margin $600,000
I = H - E Operating Income $120,000
J = H/I Degree of Operating Leverage 5 Answer #3